OpenStudy (greenpuppyz):

Reaserchers Counted 8,912 monarch butterflies on one branch of a tree at a site in Mexico.They estimated that the total number of butterflies at the site was 1,000 times as large.About how many butterflies were at the site in all? Explain your thinking,and include a statement of the solution. I think it is 8,912x1,000 on a place value chart, Is that what i am supposed to do??

OpenStudy (greenpuppyz):

I have figured out how to do it but I am not sure to divide or multiply

OpenStudy (mathstudent55):

Yes, the multiplication you show is correct. If 1 branch has 8,912 butterflies, and the estimate is that the total number of butterflies is 1000 times this amount, then there are 1,000 * 8,912 butterflies altogether. To multiply by 1,000, just move the decimal point 3 places to the right. For example, 1,000 * 5 = 5,000 1,000 * 75 = 75,000 1,000 * 2,000 = 2,000,000 Now try 1,000 * 8,912